# Robustness in Large Language Models: A Survey of Mitigation Strategies and Evaluation Metrics

### v1

@@ 5.1.5 Fairness Metrics under Stress
  • Concept: Standard fairness evaluation often checks for performance disparities or stereotypical associations on clean data. Fairness robustness evaluation applies similar principles but within a stress-testing context, using benchmarks like FLEX (Jung et al., 2025).
  • Specific Metrics: These often involve measuring differences in other robustness metrics (like ASR, refusal rates, toxicity scores, or task accuracy) across different demographic groups when the model is prompted with bias-inducing adversarial inputs (Jung et al., 2025). For example, one might measure if a jailbreak attack is more successful (higher ASR) when the harmful instruction targets a specific demographic group, or if the model refuses benign requests more often for certain groups under adversarial pressure. Metrics derived from causal frameworks, like stratified invariance, can also quantify the model’s reliance on protected attributes under intervention (Cotta & Maddison, 2024).
@@ 5.1.6 Task-Specific Robustness Metrics
  • Code Generation: Beyond functional correctness, robustness evaluation focuses on aspects like security vulnerabilities or handling of edge cases too (Rahman & Kundu, 2024). Metrics might involve:
    • Static Analysis: Checking for the presence or absence of necessary robustness features like input validation, exception handling, or specific control structures compared to human-written reference code (Li et al., 2025d).
    • Robustness to Prompt Perturbation: Evaluating if code generation remains correct when the natural language instruction contains typos or semantic variations (Gan et al., 2024).
    • Code Hallucination Metrics: Specific metrics from benchmarks like HalluCode, such as Accuracy of Hallucination Existence Recognition (Acc-Rec) (did the model correctly identify if generated code contains hallucination ?) and Accuracy of Hallucination Mitigation (Acc-Mit) (did the model successfully correct the hallucinated code ?) (Liu et al., 2024a).
  • Logical Reasoning: Evaluating performance degradation on standard reasoning benchmarks (e.g., ReClor, LogiQA) when the task structure is perturbed, for example, by shuffling the order of multiple-choice options or replacing the correct option with “none of the above” (as done in ReClor-plus, LogiQA-plus) (Bao et al., 2025). Robustness can also be tested against typographical errors in the problem statement using benchmarks like R2ATA (Gan et al., 2024).
@@ 5.1.7 Hallucination and Factuality Metrics under Stress
This crucial category focuses on measuring an LLM’s propensity to generate content that is factually incorrect, inconsistent with provided sources (unfaithful), or nonsensical (hallucination) (Liu et al., 2024a). The evaluation is particularly concerned with how this tendency changes under stressful conditions like OOD inputs, adversarial prompts, or questions probing the model’s knowledge boundaries.
@@
    • Knowledge Base (KB) / Corpus Comparison: Generated statements are checked against external structured KBs or large factual corpora (Muhlgay et al., 2024). The FACTOR benchmark operationalizes this by creating contrastive (true/false) statements from a corpus and evaluating the LM’s likelihood score assigned to the true version versus false alternatives (Muhlgay et al., 2024).
    • Exact Match (EM) (Wang et al., 2024c): Measures the percentage of generated answers that perfectly match a known factual reference answer. It is simple but often too strict, failing to credit semantically correct paraphrases.
    • Atomic Fact Verification (e.g., FactScore, FCH Rate): This approach decomposes a longer generated text into individual factual claims (“atomic facts”) and verifies each claim against a reliable source like Wikipedia (Wang et al., 2024c). The final score is often the percentage of verified facts. This allows for fine-grained analysis of long-form generation. The Factual Contradiction Hallucination (FCH) rate used in the DefAn benchmark follows a similar principle (Rahman et al., 2024).
    • FEWL (Factualness Evaluations via Weighting LLMs): A reference-free metric designed for scenarios without gold-standard answers (Wei et al., 2024). It uses multiple external LLMs (“reference LLMs”) as proxies. It calculates a truthfulness score by weighting the agreement between the evaluated response and each reference LLM’s response based on the reference LLM’s estimated “expertise” (disagreement with wrong answers) and “non-laziness” (consistency across similar questions) (Wei et al., 2024).
    • QA/NLI-based Metrics: These leverage auxiliary models. A Question Generation/Answering system can be used to generate questions from the generated text and check if the answers are consistent with the source document. Alternatively, Natural Language Inference (NLI) models can classify the relationship (entailment, contradiction, neutral) between generated statements and source text sentences (Chen et al., 2024).
@@
    • Methods often overlap with fact-checking (e.g., using NLI or QA models) but focus specifically on adherence to the provided context rather than general world knowledge (Chen et al., 2024). Faithfulness metrics evaluate if all claims in the output can be inferred from the source (Yang et al., 2025a).
    • Mol-Hallu: A domain-specific metric for evaluating faithfulness in molecular comprehension tasks. It uses an entailment model to check if scientific entities mentioned in the generated text are actually supported by the provided molecular descriptions (Li et al., 2025a).
@@
    • Self-Consistency Checks (e.g., SelfCheckGPT): This approach involves generating multiple responses from the same LLM for a single prompt, typically using temperature sampling (T>0) to introduce diversity (Wei et al., 2024). The consistency (or lack thereof) across these responses is then measured. High variance or contradiction among samples might indicate hallucination, while consistent outputs suggest higher confidence or factuality. Various metrics can be used to quantify this, such as semantic similarity variance or contradiction detection using NLI models (Wei et al., 2024).
    • LLM-based Consistency Score: This method uses a secondary model to evaluate consistency by generating multiple responses (with controlled randomness) to the same prompt, then calculating how often these responses logically align (Saxena et al., 2024). The resulting “consistency score” measures reliability under variable conditions.
    • MetaQA: This method applies metamorphic testing principles (Yang et al., 2025a). It generates variations of the original prompt using synonym or antonym substitutions. It then checks if the LLM’s responses to these related prompts exhibit the expected consistency (for synonyms) or inconsistency (for antonyms). Deviations from expected patterns can signal factual errors or hallucinations (Yang et al., 2025a).
@@
    • Confidence Scores: Using metrics derived from the model’s output probabilities, such as token-level probabilities, sequence log-probability, or entropy (Mahaut et al., 2024). The assumption is that lower confidence might correlate with a higher likelihood of hallucination. However, this requires the model to be well-calibrated for the confidence scores to be reliable indicators.
    • Energy Score: While primarily used for OOD detection, the energy score calculated for reward models can indicate when the model is operating on inputs (prompts or responses) far from its training distribution, suggesting its reward estimate (which might relate to factuality) could be unreliable (LeVine et al., 2024).
@@ 5.2 Benchmarks
Complementing the metrics described above, a growing number of benchmarks and evaluation frameworks are being developed to systematically assess LLM robustness across its various dimensions. These benchmarks provide standardized datasets and protocols for testing models under specific types of stress.
@@
AdvGLUE++ MNLI, QQP, etc. Perturbed versions of GLUE using textual attacks
@@
R2ATA Adversarial Resilience (Reasoning), Noise Perturbed GSM8K, BBH, MMLU Adversarial typographical errors in reasoning problems
@@
5.2.1 Adversarial Benchmarks
  • AdvGLUE / AdvGLUE++: These benchmarks are adversarial versions of the popular GLUE (General Language Understanding Evaluation) benchmark. They apply various textual adversarial attack methods (e.g., word substitutions using synonyms, character-level perturbations, sentence paraphrasing) to the original GLUE tasks (like sentiment analysis, natural language inference). AdvGLUE++ (Wang et al., 2021) expands upon AdvGLUE with more adversarial examples, targeting newer LLMs. They serve as a standard for evaluating the robustness of fundamental language understanding capabilities against common textual attacks.
  • PromptBench: Unlike benchmarks focusing on perturbing input samples, PromptBench (Zhang, 2023) specifically evaluates the robustness of LLMs to adversarial prompts (the instructions given to the model). It includes a framework for dynamically generating adversarial prompts using character, word, sentence, and semantic-level attacks and applies these perturbed prompts across various tasks (sentiment analysis, NLI, QA, translation, math) and datasets. This targets the sensitivity of LLMs to the way instructions are phrased.
  • Jailbreak Benchmarks (e.g., JailbreakBench, AdvBench subset): These benchmarks consist of prompts specifically designed to circumvent an LLM’s safety alignment training and elicit responses that are harmful, unethical, illegal, or otherwise prohibited by the model’s safety guidelines (Xhonneux et al., 2024). Examples include prompts using role-playing scenarios, hypothetical situations, or character encodings to trick the model. Evaluating performance on these benchmarks (often using ASR or Compliance Ratio) is crucial for assessing the effectiveness of safety measures.
  • R2ATA (Gan et al., 2024): This benchmark focuses on the robustness of LLM reasoning capabilities. It uses the Adversarial Typo Attack (ATA) algorithm to introduce minimal, targeted typographical errors into reasoning problems (from datasets like GSM8K, BBH, MMLU). It then measures the degradation in the LLM’s ability to perform correct step-by-step reasoning (Chain-of-Thought) and arrive at the right answer.
Adversarial benchmarks, while primarily focused on inducing misclassification or safety failures, can inadvertently trigger hallucinations. Attacks that manipulate reasoning processes or push models into unusual states via jailbreaking prompts (Gan et al., 2024; Zhou et al., 2024a) might cause them to generate factually incorrect or nonsensical outputs as a byproduct of the attack’s success or the model’s attempt to handle the confusing input. Evaluating factuality metrics on adversarial datasets can provide insights into how robust a model’s factual grounding is when under direct attack.
@@ 5.2.2 OOD Benchmarks
  • BOSS (Benchmark suite for Out-of-distribution robustneSS): BOSS 8 is a notable effort to create a more challenging and holistic OOD evaluation suite for NLP. It covers five diverse tasks: Sentiment Analysis (SA), Toxic Detection (TD), Natural Language Inference (NLI), Named Entity Recognition (NER), and Extractive Question Answering (EQA) (Yuan et al., 2023). For each task, BOSS includes one large, diverse In-Distribution (ID) dataset (e.g., Amazon reviews for SA, MNLI for NLI, SQuAD for EQA) and three corresponding OOD datasets (Yuan et al., 2023). The OOD datasets are carefully selected based on a protocol that prioritizes: (1) distinct data sources/domains to ensure low semantic similarity (measured via SimCSE) with the ID set and (2) evidence of a significant performance drop for a baseline model trained on the ID set, indicating a challenging distribution shift (Yuan et al., 2023). Examples of OOD datasets included are Dynasent, SemEval (SA); ANLI, ContractNLI, WANLI (NLI); AdvQA, NewsQA, SearchQA (EQA) (Yuan et al., 2023).
  • ANLI (Adversarial NLI): Created through an iterative human-and-model-in-the-loop process, ANLI contains NLI examples specifically designed to be difficult for contemporary models. Due to its challenging nature and distinct creation process, it is frequently used as an OOD benchmark for NLI task generalization (Yuan et al., 2023).
  • Dynasent: A sentiment analysis dataset generated dynamically using adversarial methods to create challenging examples that models trained on standard sentiment datasets might misclassify (Yuan et al., 2023).
  • WANLI: An NLI dataset synthesized using GPT-3, focusing on challenging linguistic patterns that models often struggle with in datasets like MNLI (Yuan et al., 2023).
  • OODRobustBench (Li et al., 2024a): This is a comprehensive benchmark to evaluate adversarial robustness under dataset and threat distribution shifts. Further, this benchmark demonstrates that adversarial robustness suffers significant degradation under distribution shifts, yet in-distribution robustness correlates strongly and linearly with out-of-distribution robustness.

### v2

@@ 5.1.5 Fairness Metrics under Stress
  • Concept: Standard fairness evaluation often checks for performance disparities or stereotypical associations on clean data. Fairness robustness evaluation applies similar principles but within a stress-testing context, using benchmarks like FLEX (jung2025flexbenchmarkevaluatingrobustness).
  • Specific Metrics: These often involve measuring differences in other robustness metrics (like ASR, refusal rates, toxicity scores, or task accuracy) across different demographic groups when the model is prompted with bias-inducing adversarial inputs (jung2025flexbenchmarkevaluatingrobustness). For example, one might measure if a jailbreak attack is more successful (higher ASR) when the harmful instruction targets a specific demographic group, or if the model refuses benign requests more often for certain groups under adversarial pressure. Metrics derived from causal frameworks, like stratified invariance, can also quantify the model’s reliance on protected attributes under intervention (cotta2024testtimefairnessrobustnesslarge).
@@ 5.1.6 Task-Specific Robustness Metrics
  • Code Generation: Beyond functional correctness, robustness evaluation focuses on aspects like security vulnerabilities or handling of edge cases too (Rahman2024CodeH). Metrics might involve:
    • Static Analysis: Checking for the presence or absence of necessary robustness features like input validation, exception handling, or specific control structures compared to human-written reference code (unknowntask).
    • Robustness to Prompt Perturbation: Evaluating if code generation remains correct when the natural language instruction contains typos or semantic variations (gan2024reasoningrobustnessllmsadversarial).
    • Code Hallucination Metrics: Specific metrics from benchmarks like HalluCode, such as Accuracy of Hallucination Existence Recognition (Acc-Rec) (did the model correctly identify if generated code contains hallucination ?) and Accuracy of Hallucination Mitigation (Acc-Mit) (did the model successfully correct the hallucinated code ?) (liu2024exploringevaluatinghallucinationsllmpowered).
  • Logical Reasoning: Evaluating performance degradation on standard reasoning benchmarks (e.g., ReClor, LogiQA) when the task structure is perturbed, for example, by shuffling the order of multiple-choice options or replacing the correct option with “none of the above” (as done in ReClor-plus, LogiQA-plus) (bao2025assessingenhancingrobustnesslarge). Robustness can also be tested against typographical errors in the problem statement using benchmarks like R2ATA (gan2024reasoningrobustnessllmsadversarial).
@@ 5.1.7 Hallucination and Factuality Metrics under Stress
This crucial category focuses on measuring an LLM’s propensity to generate content that is factually incorrect, inconsistent with provided sources (unfaithful), or nonsensical (hallucination) (liu2024exploringevaluatinghallucinationsllmpowered). The evaluation is particularly concerned with how this tendency changes under stressful conditions like OOD inputs, adversarial prompts, or questions probing the model’s knowledge boundaries.
@@
    • Knowledge Base (KB) / Corpus Comparison: Generated statements are checked against external structured KBs or large factual corpora (muhlgay2024generatingbenchmarksfactualityevaluation). The FACTOR benchmark operationalizes this by creating contrastive (true/false) statements from a corpus and evaluating the LM’s likelihood score assigned to the true version versus false alternatives (muhlgay2024generatingbenchmarksfactualityevaluation).
    • Exact Match (EM) (wang2024factualitylargelanguagemodels): Measures the percentage of generated answers that perfectly match a known factual reference answer. It is simple but often too strict, failing to credit semantically correct paraphrases.
    • Atomic Fact Verification (e.g., FactScore, FCH Rate): This approach decomposes a longer generated text into individual factual claims (“atomic facts”) and verifies each claim against a reliable source like Wikipedia (wang2024factualitylargelanguagemodels). The final score is often the percentage of verified facts. This allows for fine-grained analysis of long-form generation. The Factual Contradiction Hallucination (FCH) rate used in the DefAn benchmark follows a similar principle (rahman2024defandefinitiveanswerdataset).
    • FEWL (Factualness Evaluations via Weighting LLMs): A reference-free metric designed for scenarios without gold-standard answers (wei2024measuringreducingllmhallucination). It uses multiple external LLMs (“reference LLMs”) as proxies. It calculates a truthfulness score by weighting the agreement between the evaluated response and each reference LLM’s response based on the reference LLM’s estimated “expertise” (disagreement with wrong answers) and “non-laziness” (consistency across similar questions) (wei2024measuringreducingllmhallucination).
    • QA/NLI-based Metrics: These leverage auxiliary models. A Question Generation/Answering system can be used to generate questions from the generated text and check if the answers are consistent with the source document. Alternatively, NLI models can classify the relationship (entailment, contradiction, neutral) between generated statements and source text sentences (chen2024hallucinationdetectionrobustlydiscerning).
@@
    • Methods often overlap with fact-checking (e.g., using NLI or QA models) but focus specifically on adherence to the provided context rather than general world knowledge (chen2024hallucinationdetectionrobustlydiscerning). Faithfulness metrics evaluate if all claims in the output can be inferred from the source (yang2025hallucinationdetectionlargelanguage).
    • Mol-Hallu: A domain-specific metric for evaluating faithfulness in molecular comprehension tasks. It uses an entailment model to check if scientific entities mentioned in the generated text are actually supported by the provided molecular descriptions (li2025detectdefeatmolecularmirage).
    LLMs
    Accuracy/ F1 Drop (§ 5.1.1) If Drop > X pp 1. Add domain data 2. Data Augmentation 3. Distributionally Robust Optimization (DRO) 4. Test-time adaptation Attack Success Rate (ASR) & Compliance Rate (§ 5.1.1) If ASR > Y % 1. Adversarial fine-tuning 2. RLHF/DPO safety 3. Integrate robust prompting templates and constrained decoding 4. Output filters 5. Ensembling judges OOD Performance (F1, Gap, AUROC) (§ 5.1.2) If OOD Gap is large or AUROC < 0.8 1. DRO Training 2. Collect OOD-aware data 3. Deploy OOD Detectors 4. RAG for grounding Prompt Consistency (CR, SC, RC) (§ 5.1.3) If CR < 90% 1. Paraphrase Augmentation 2. Consistency Regularization 3. Prompt Canonicalization Calibration (ECE, Brier Score) (§ 5.1.4) If ECE is high 1. Temperature Scaling 2. Calibration-aware losses 3. Model ensembles or Bayesian approximates Fairness under Stress (FLEX) (§ 5.1.5) If disparity is high under stress prompts 1. Debias or reweight datasets 2. Fairness-aware tuning 3. Post-processing filters Hallucination / Factuality (HRI, FactScore, FCH) (§ 5.1.7) If HRI is high or FactScore is low 1. Adopt RAGs with trusted knowledge sources 2. Fact-checking layers 3. Grounded QA fine-tuning
    Figure 6: Actionable guidance for LLM practitioners where performance metrics can be taken from Table 4, and by identifying the threat model or corresponding weak dimensions, practitioners can apply corresponding strategies from Table 3.
@@
    • Self-Consistency Checks (e.g., SelfCheckGPT): This approach involves generating multiple responses from the same LLM for a single prompt, typically using temperature sampling (T>0) to introduce diversity (wei2024measuringreducingllmhallucination). The consistency (or lack thereof) across these responses is then measured. High variance or contradiction among samples might indicate hallucination, while consistent outputs suggest higher confidence or factuality. Various metrics can be used to quantify this, such as semantic similarity variance or contradiction detection using NLI models (wei2024measuringreducingllmhallucination).
    • LLM-based Consistency Score: This method uses a secondary model to evaluate consistency by generating multiple responses (with controlled randomness) to the same prompt, then calculating how often these responses logically align (saxena2024evaluatingconsistencyreasoningcapabilities). The resulting “consistency score” measures reliability under variable conditions.
    • MetaQA: This method applies metamorphic testing principles (yang2025hallucinationdetectionlargelanguage). It generates variations of the original prompt using synonym or antonym substitutions. It then checks if the LLM’s responses to these related prompts exhibit the expected consistency (for synonyms) or inconsistency (for antonyms). Deviations from expected patterns can signal factual errors or hallucinations (yang2025hallucinationdetectionlargelanguage).
@@
    • Confidence Scores: Using metrics derived from the model’s output probabilities, such as token-level probabilities, sequence log-probability, or entropy (Mahaut_2024). The assumption is that lower confidence might correlate with a higher likelihood of hallucination. However, this requires the model to be well-calibrated for the confidence scores to be reliable indicators.
    • Energy Score: While primarily used for OOD detection, the energy score calculated for reward models can indicate when the model is operating on inputs (prompts or responses) far from its training distribution, suggesting its reward estimate (which might relate to factuality) could be unreliable (levine2024baselineanalysisrewardmodels).
@@ 5.2 Benchmarks
Complementing the metrics described above, a growing number of benchmarks and evaluation frameworks are being developed to systematically assess LLM robustness across its various dimensions. Table 5 provides an overview of the benchmarks discussed. These benchmarks provide standardized datasets and protocols for testing models under specific types of stress.
5.2.1 Adversarial Benchmarks
  • AdvGLUE / AdvGLUE++: These benchmarks are adversarial versions of the popular GLUE (General Language Understanding Evaluation) benchmark. They apply various textual adversarial attack methods (e.g., word substitutions using synonyms, character-level perturbations, sentence paraphrasing) to the original GLUE tasks (like sentiment analysis, NLI). AdvGLUE++ (Wang2021AdversarialGA) expands upon AdvGLUE with more adversarial examples, targeting newer LLMs. They serve as a standard for evaluating the robustness of fundamental language understanding capabilities against common textual attacks.
  • PromptBench: Unlike benchmarks focusing on perturbing input samples, PromptBench (zhang2023awesome) specifically evaluates the robustness of LLMs to adversarial prompts (the instructions given to the model). It includes a framework for dynamically generating adversarial prompts using character, word, sentence, and semantic-level attacks and applies these perturbed prompts across various tasks (sentiment analysis, NLI, QA, translation, math) and datasets. This targets the sensitivity of LLMs to the way instructions are phrased.
  • Jailbreak Benchmarks (e.g., JailbreakBench, AdvBench subset): These benchmarks consist of prompts specifically designed to circumvent an LLM’s safety alignment training and elicit responses that are harmful, unethical, illegal, or otherwise prohibited by the model’s safety guidelines (xhonneux2024efficientadversarialtrainingllms). Examples include prompts using role-playing scenarios, hypothetical situations, or character encodings to trick the model. Evaluating performance on these benchmarks (often using ASR or Compliance Ratio) is crucial for assessing the effectiveness of safety measures.
  • R2ATA (gan2024reasoningrobustnessllmsadversarial): This benchmark focuses on the robustness of LLM reasoning capabilities. It uses the Adversarial Typo Attack (ATA) algorithm to introduce minimal, targeted typographical errors into reasoning problems (from datasets like GSM8K, BBH, MMLU). It then measures the degradation in the LLM’s ability to perform correct step-by-step reasoning (Chain-of-Thought) and arrive at the right answer.
Adversarial benchmarks, while primarily focused on inducing misclassification or safety failures, can inadvertently trigger hallucinations. Attacks that manipulate reasoning processes or push models into unusual states via jailbreaking prompts (gan2024reasoningrobustnessllmsadversarial; zhou2024robustpromptoptimizationdefending) might cause them to generate factually incorrect or nonsensical outputs as a byproduct of the attack’s success or the model’s attempt to handle the confusing input. Evaluating factuality metrics on adversarial datasets can provide insights into how robust a model’s factual grounding is when under direct attack.
@@ 5.2.1 Adversarial Benchmarks
AdvGLUE++ MNLI, QQP, etc.
@@
Adversarial Resilience (Reasoning), Noise R2ATA Perturbed GSM8K, BBH, MMLU Adversarial typographical errors in reasoning problems
@@ 5.2.2 OOD Benchmarks
  • BOSS (Benchmark suite for Out-of-distribution robustneSS): BOSS is a notable effort to create a more challenging and holistic OOD evaluation suite for NLP. It covers five diverse tasks: Sentiment Analysis (SA), Toxic Detection (TD), NLI, Named Entity Recognition (NER), and Extractive Question Answering (EQA) (yuan2023revisiting). For each task, BOSS includes one large, diverse In-Distribution (ID) dataset (e.g., Amazon reviews for SA, MNLI for NLI, SQuAD for EQA) and three corresponding OOD datasets (yuan2023revisiting). The OOD datasets are carefully selected based on a protocol that prioritizes: (1) distinct data sources/domains to ensure low semantic similarity (measured via SimCSE) with the ID set and (2) evidence of a significant performance drop for a baseline model trained on the ID set, indicating a challenging distribution shift (yuan2023revisiting). Examples of OOD datasets included are Dynasent, SemEval (SA); ANLI, ContractNLI, WANLI (NLI); AdvQA, NewsQA, SearchQA (EQA) (yuan2023revisiting).
  • ANLI (Adversarial NLI): Created through an iterative human-and-model-in-the-loop process, ANLI contains NLI examples specifically designed to be difficult for contemporary models. Due to its challenging nature and distinct creation process, it is frequently used as an OOD benchmark for NLI task generalization (yuan2023revisiting).
  • Dynasent: A sentiment analysis dataset generated dynamically using adversarial methods to create challenging examples that models trained on standard sentiment datasets might misclassify (yuan2023revisiting).
  • WANLI: An NLI dataset synthesized using GPT-3, focusing on challenging linguistic patterns that models often struggle with in datasets like MNLI (yuan2023revisiting).
  • OODRobustBench (li2024oodrobustbench): This is a comprehensive benchmark to evaluate adversarial robustness under dataset and threat distribution shifts. Further, this benchmark demonstrates that adversarial robustness suffers significant degradation under distribution shifts, yet in-distribution robustness correlates strongly and linearly with out-of-distribution robustness.